OpenStudy (anonymous):

. If you want to save $25,000 for a down payment on a house and you have ten years to save this amount, how much would you need to save monthly to achieve this goal if the interest rate is 5% compounded monthly. This was my work, but Im not sure if it is right.. A=p(1=r/n)^nt 1/p=(1+0.05)^12*10/25000 1/p=(1.0416)^120/25000 1/p=(133.081151)/25000 P*133.081151 = 250000 250000/133.081151= $187.86

OpenStudy (anonymous):

Follow this formula: 1/P= (1 + r/n)^nt /A P= what you deposit monthly. R= rate of interest. T= number of years deposited. A= amount of money accumulated. N= number of times the interest is compounded per year

OpenStudy (anonymous):

can you try to do it?

OpenStudy (anonymous):

1/P= (1 + .05/12)^12 x 10 /25,000 1/P= (1.0416)^120 /25,000 1/P=(133.081151)/25,000 P x 133.081151 = 25,000 (Cross multiplied) P = 187.86

OpenStudy (anonymous):

it correct then ^^

OpenStudy (tkhunny):

It is very, very difficult to follow what you are trying to do. "A=p(1+r/n)^nt" This is for only one payment. How does this manage to get to look like something for more than one payment? Try an accumulation of MULTIPLE periodic payments at the beginnnig of each period! i = 0.05 j = 0.05/12 r = 1+j \(25000 = Pmt\left(\dfrac{r^{120}-1}{r-1}\right)\cdot r\)

OpenStudy (anonymous):

To check, plug into the original equation. (Before I changed it to solve to P) A= P(1 + r/n)^nt 25000= 187.86 (1 + .05/12) ^12 x 10 25000= 187.86 (1.0416)^120 25000= 187.86 (133.08115) 25000= 25000

OpenStudy (anonymous):

so it is correct your answer it should be like that

OpenStudy (tkhunny):

Pmt = 160.33 Until you abandon that single payment structure, you will not find the correct result.

OpenStudy (anonymous):

you can check with 160.33 you should find same answer

OpenStudy (anonymous):

it is not geometric series it is compound interest

OpenStudy (tkhunny):

It is a compound interest problem that can be solved by the realization that it is represented perfectly with a geometric series. The distinction is unnecessary.

OpenStudy (anonymous):

Thank you.